OpenStudy (anonymous):

The ratio of male students to female students at a certain university is 5 to 4. If there is a total of 6966 students, find the number of male students and the number of female students.

OpenStudy (anonymous):

6966/(5+4) = 774 774*5 = 3870 male students 774*4 = 3096 female students